OpenStudy (anonymous):

Multiply (c+(1)/(3))(c+(1)/(4))

OpenStudy (anonymous):

I got this for my answer am I correct? c^2+17c/12+1/12

OpenStudy (anonymous):

No.\[(c+(1)/(3))(c+(1)/(4)) \]\[\left(c+\frac{1}{3}\right)\left(c+\frac{1}{4}\right) \]\[c^2+\frac{7 c}{12}+\frac{1}{12} \]